How Many 8inch x 8 feet wood boards are needed to fill up an area that equals 1000 square feet?
Answer by swincher4391(1107)

8 inch = 2/3 feet
area of one board = 8 feet * 2/3 feet = 16/3 feet^2
1000/(16/3) = 187.5 boards.
We can't have half of a board (well maybe you can), so we only use 187 full boards.